You manage a pension fund, and your liabilities consist of three payments as follows:

Time Payment
5 years 5 million
10 years 10 million
30 years 30 million

The term structure is currently flat at 10%

a) Compute the present value of your liabilities.

b) Compute modified duration of your liabilities.
Answer #1

a.

Present Value = 5/(1.10)5 + 10/(1.10)10 + 30/(1.10)30

Present Value = $8.68 million

b.

Macaulay Duration = [(5)5/(1.10)5 + (10)10/(1.10)10 + (30)30/(1.10)30](8.68)

Macaulay Duration = (15.52 + 38.55 + 51.58)/8.68

Macaulay Duration = 12.17 years

Modified Duration = 12.17/(1.10)

Modified Duration = 11.07 years